OpenStudy (kobeni-chan):

Will medal and fan!! Find all solutions to the equation. (cos^2)x + (2 cos x) + 1 = 0

OpenStudy (kobeni-chan):

@Michele_Laino

hartnn (hartnn):

could you first find 2 values of cos x ? its a quadratic in cos x......

hartnn (hartnn):

let y = cos x y^2 +2y +1 = 0

OpenStudy (kobeni-chan):

Sorry I'm confused >_< From what they've been teaching me, I'm supposed to isolate the trig. part, and then add 2πn to each solution..?

hartnn (hartnn):

yes, all that was just to isolate (or find the multiple values of) the trig part which is cos x here. since the equation is quadratic in cos x, it can have multiple values. hence first find them, then we can proceed.

hartnn (hartnn):

let y = cos x y^2 +2y +1 = 0 y =... hence, cos x = ...

OpenStudy (kobeni-chan):

Ohh ok. Is it +1 and +1 again? :/ or (y+1)^2

hartnn (hartnn):

(y+1)^2 = 0 so y+1 = 0 so, y = -1 cos x = -1 for what values does cos give you -1?